40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jan 27 14:40:19 2022, shruti, Update, Noise Budget, Measurement and calibration of noise spectra Spectra20220127.pdfCalibratedNoiseSpectra20220127.pdfCavityMeasurements.pdf
    Reply  Wed Feb 2 11:08:47 2022, shruti, Update, Noise Budget, Measurement and calibration of noise spectra OLTF_20220201.pdfCalibratedNoiseSpectra.pdfCalibratedNoiseSpectra20220127_corrected.pdfattn10dB.jpegFilterTF.pdf
       Reply  Mon Feb 7 11:55:07 2022, shruti, Update, Noise Budget, Updated pomona filter, loop model filter.jpegFilterTF.pdfLoopModel.pdf
          Reply  Wed Feb 9 16:25:54 2022, rana, Update, Noise Budget, Updated pomona filter, loop model 
             Reply  Tue Feb 15 11:21:18 2022, shruti, Update, Noise Budget, Updated loop model LoopModel.pdfLB_meas_setup.jpeg
                Reply  Tue Feb 15 16:56:50 2022, rana, Update, Noise Budget, Updated loop model 
Message ID: 2846     Entry time: Thu Jan 27 14:40:19 2022     Reply to this: 2850
Author: shruti 
Type: Update 
Category: Noise Budget 
Subject: Measurement and calibration of noise spectra 

Calibrating the noise budget

  1. I measured the noise spectra at the control point (after the servo) and the error point (before the servo and input attenuator) using the Moku. See these (loop, setup) diagrams for a better description. For the previous measurement (elog 2775 and elog 2776) we had chosen the ambiguous units of dBm/Hz, but this time we used the noise in dBV/rt Hz which does not require the impedance. I took separate measurements for the different frequency regions and stitched the data together. The raw noise spectra at both these points are shown in Attachment 1.
  2. Using the PDH signal on the oscilloscope I estimated the cavity pole frequency using the same process outlined in the table here. I found the cavity pole frequency to be 2.3 MHz and the cavity response to be 132 mV/Hz -- the two quantities required to estimate the cavity transfer function (modeled as a single pole low pass filter). [update 27-Jan-22: Added Attachment 3 with the crude cavity pole and response measurements]
  3. I also used the open loop transfer function measured earlier to correct for the loop suppression
  4. I used the updated noise calibration notebook that I uploaded here and the data (uploaded here).
  5. Attachment 2 shows a calibrated noise spectra using only the noise measured at the PDH error point. The red curves is the measured noise in-loop calibrated to the frequency noise at cavity input, the blue curves incorporate the inverted loop suppression factor. The lighter curves show the previously measured noise spectra without the cantilever.

All this was measured without adding the pomona box filter which would be my next step.

 

Attachment 1: Spectra20220127.pdf  72 kB  Uploaded Fri Jan 28 16:01:33 2022  | Hide | Hide all
Spectra20220127.pdf
Attachment 2: CalibratedNoiseSpectra20220127.pdf  129 kB  Uploaded Fri Jan 28 16:40:53 2022  | Hide | Hide all
CalibratedNoiseSpectra20220127.pdf
Attachment 3: CavityMeasurements.pdf  636 kB  Uploaded Mon Jan 31 11:18:44 2022  | Hide | Hide all
CavityMeasurements.pdf
ELOG V3.1.3-